A Pareto distribution is often used in economics to explain a distribution of wealth. Let a random variable X have a Pareto distribution with parameter θ so that its probability distribution function is

fx(2:0) = A.c-4

for x Στο and 0 otherwise. The parameters $\theta>1$ and 0 < 0x are known and fixed; $A$ is a constant to be determined.

a) Assuming that $\theta>3,$ find the expected value and variance of $X$ ?

b) Show that for 3 ≥ θ > 2 the Pareto distribution has a finite mean but infinite variance, and for 2 ≥ θ > 1 both mean and variance are infinite.
Answer #1

(a) E(X) = A xrx-0 dx - 4 xr-O+1 % (r-o +1) Ixo E(X) = A x-0 0 -o (9 ) - Axo 0 (0-1) xo = (0-1) E (x2)= A x (0-1), co -(0-1)1<O<2 E(x) = x so v(x) - x 1. < 0-2 < 0 it 022 as E(X) - 0.8 AXO dx = 1 x-0+1 10 -oti = 1 I Xo - A xo-(0-1) X (0-1) (0-1) & A (0-1) X